User Tools

Site Tools


lbaops:lbajul2011:v440atilog

Recording to ATNF V011/B1-B7 mounted as /data/xraid1

Sky frequenfies 24975 MHz (later changed to 24975 MHz) for LCP/RCP, 2nd LO 481.4 MHz

Observers: shinji horiuch and Team-B/D operation teams

08:00:00 started on time

08:37:00 DAS Frequency changed to 24967 MHz per urgently requested by PI.

08:37:00-08:42:20 No IF power due to accidental LNA amp down.

RTFC found no fringes at 12:48:40, but fringes were present again at 13:52:50

(Note: Tid was slewing from 12:40:00 to 12:54:08 to get 1921-231 on point, then slewing from 12:58:00 to 13:48:03 to 351.78-0)

14:00:00 end of Tid schedule

Note added on 5 Oct 2011: Log files have been uploaded to ftp incoming area. The tsys files are powermeter outputs scaled with a plausible Tsys value, and error may be up to 20-40 percent. Hope grobal self-cal can correct the error.

lbaops/lbajul2011/v440atilog.txt · Last modified: 2015/12/18 16:38 (external edit)